### v1

# A Generalized Analytical Framework for the Nonlinear Best-Worst Method

Abstract
To eliminate the need for optimization software in calculating weights using the nonlinear model of the Best-Worst Method (BWM), Wu et al. proposed an analytical framework for deriving optimal interval-weights. They also introduced a secondary objective function to select the best optimal weight set. However, their framework is only compatible with a single Decision-Maker (DM) and preferences quantified using the Saaty scale. In this research, we generalize their framework to accommodate any number of DMs and any scale. We first derive an analytical expression for optimal interval-weights and select the best optimal weight set. After demonstrating that the values of Consistency Index (CI) for the Saaty scale in the existing literature are not well-defined, we derive n formula for computing CI. We also derive analytical expressions for the Consistency Ratio (CR), enabling its use as an input-based consistency indicator and proving that CR satisfies some key properties, ensuring its reliability as a consistency indicator. Furthermore, we observe that criteria with equal preferences may get different weights when multiple best/worst criteria are present. To address this limitation, we modify the original optimization model for weight computation in such instances, solve it analytically to obtain optimal interval-weights, and select the best optimal weight set. Finally, we demonstrate and validate the proposed approach using numerical examples.
Keywords: Multi-criteria decision-making, Best-worst method, Optimal weights, Consistency index, Consistency ratio
1 Introduction
Multi-Criteria Decision-Making (MCDM), a fundamental branch of operations research, addresses complex decision scenarios involving conflicting criteria. Typically, MCDM approaches solve decision problems through two sequential steps: determining weights of decision criteria and ranking of alternatives. This classification divides MCDM methods into two categories: weight determination methods (such as AHP [37], ANP [38] and SMARTS [15]) and ranking methods (including TOPSIS [20], VIKOR [29], ELECTRE [36] and PROMETHEE [5]).
The Best-Worst Method (BWM) is a widely used weight calculation method that utilizes pairwise comparisons between decision criteria to derive weights [32]. In this method, the optimal weights are obtained by minimizing the distance between weight ratios and given comparison values. Based on different distance functions, various models of BWM have been developed. The original model of BWM, proposed by Rezaei [32], uses maximum deviation as the distance function. Since this approach involves solving a nonlinear optimization problem, it is known as the nonlinear BWM. Kocak et al. [22] introduced a model of BWM based on Euclidean distance, termed the Euclidean BWM. Brunelli and Rezaei [8] proposed the multiplicative BWM, incorporating the metric $\max\{\nicefrac{{x}}{{y}},\nicefrac{{y}}{{x}}\}$ defined on $(\mathbb{R_{+}},\cdot,\leq)$ into the BWM framework. Amiri and Emamat [3] further developed a goal programming-based BWM using the taxicab distance. Tu et al. [42] introduced two alternative BWM formulations: the approximate eigenvalue model and the logarithmic least squares model. For group decision-making, Safarzadeh et al. [39] proposed two extensions, one based on total deviation and the other based on maximum deviation. Xu and Wang [48] developed eleven models for individual DMs and nine for group decision-making contexts. Mohammadi and Rezaei [28] incorporated Bayesian into the BWM framework, establishing a probabilistic approach to group decision-making. Corrente et al. [11] developed the parsimonious BWM, an enhanced version of the nonlinear BWM specifically designed for decision contexts involving numerous alternatives.
To address the non-uniqueness of optimal weight sets in the nonlinear BWM, Rezaei [33] derived optimal interval-weights by formulating two optimization models. Later, he introduced the concentration ratio to measure the dispersion of these interval-weigths [34]. He also developed the linear BWM [33], which retains the underlying philosophy of the nonlinear BWM while transforming its optimization framework into a linear formulation. While this model guarantees a unique weight set, its feasible region differs from the original nonlinear approach. Wu et al. [47] developed an analytical approach to derive optimal interval-weights for the nonlinear BWM, eliminating the model’s dependency on optimization software. Building on this foundation, they incorporated a secondary objective function to determine the best optimal weight set from the solution space. Ratandhara and Kumar [31] subsequently proposed an analytical framework for the multiplicative BWM, achieving the same objectives of software independence and selection of the best optimal weight set.
Consistency measurement of decision data is crucial in MCDM methods since outcomes depend directly on this input. In BWM, consistency evaluation is performed through the Consistency Ratio (CR), which is typically computed using the optimal objective value and Consistency Index (CI) [32]. This output-based consistency indicator can only provide feedback about inconsistencies after completing all calculations, resulting in reduced time efficiency. Liang et al. [26] introduced an alternative input-based consistency indicator called input-based CR, along with establishing threshold values for both output-based and input-based CR to check admissibility of preference values. Furthermore, Lei et al. [24] developed an optimization model that recommends optimal preference modifications while achieving both ordinal consistency and an acceptable level of cardinal consistency.
To handle uncertain preferences, several fuzzy-set-based extensions of BWM have been proposed. Guo and Zhao [17] extended the nonlinear BWM to a fuzzy environment, while Rostami et al. [35] introduced a fuzzy adaptation of the goal programming-based BWM. Additionally, Ratandhara and Kumar [30] proposed an $\alpha$-cut interval-based model of fuzzy BWM. The BWM framework has also been extended to more advanced uncertainty representations, including intuitionistic fuzzy sets [46, 45, 10], hesitant fuzzy sets [2, 25] and spherical fuzzy sets [19]. Moreover, BWM has been integrated with other MCDM techniques, such as BWM-VIKOR [1, 16], BWM-ELECTRE [49, 9], BWM-TOPSIS [44, 41], BWM-MULTIMOORA [18, 50] and Best-Worst Tradeoff (BWT) method [27]. Owing to its simplicity and reliability, BWM has been widely applied in practical decision-making problems, notably in location selection [43, 4], logistics risk assessment [12, 13] and supplier selection [40, 44].
In this study, we establish the following key research gaps in the nonlinear BWM.
  1. (i)
    The framework proposed by Wu et al. [47] works well when preferences are quantified using the Saaty scale, but leads to different optimal interval-weights and a best optimal weight set than their actual values (in fact, optimal interval-weights are not well-defined) when some other scale such as the Salo-Hämäläinen, Lootsma or Donegan-Dodd-McMaster scale is used.
  2. (ii)
    The existing framework leads to non-well-defined optimal interval-weights and a different best weight set than the actual one, even when preferences are quantified using the Saaty scale in the presence of multiple DMs.
  3. (iii)
    The values of CI for the Saaty scale computed by Rezaei [32] are not well-defined.
  4. (iv)
    In the presence of multiple best/worst criteria, computing weights by arbitrarily selecting any one best/worst criterion leads to different weights (both interval-weights and the best weights) for criteria having equal preference.
In this work, we propose a generalized analytical framework for the nonlinear BWM that is compatible with any preference scale and any number of DMs. Our approach derives optimal interval-weights by solving an optimal modification-based optimization problem, which is analytically equivalent to the original optimization model. From the collection of all optimal weight sets, we also select the best optimal weight set. Based on this framework, we derive an exact formula for the CI valid for any scale, along with an analytical expression for the CR that serves as an input-based consistency indicator. We further establish essential properties of CR to verify its validity as a consistency measure. To ensure equal weights for criteria with equal preferences when multiple best/worst criteria exist, we develop a modified optimization model and derive corresponding optimal interval-weights and the best optimal weight set analytically. We validate the proposed approach and demonstrate its effectiveness through numerical examples.
The remainder of this paper is organized as follows. Section 2 introduces fundamental definitions and provides a concise overview of the nonlinear BWM and its existing framework. Section 3 identifies and examines critical research gaps in the current methodology. In Section 4, we present our generalized analytical framework for the nonlinear BWM, supported by demonstrative numerical examples. Finally, Section 5 discusses concluding remarks and suggests potential directions for future research.
2 Preliminary
In this section, we discuss some fundamental definitions and key results, along with a brief overview of the nonlinear BWM and its analytical framework.
2.1 Basic Concepts and Results
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of decision criteria, and let $D=\{c_{1},c_{2},\ldots,c_{n}\}\setminus\{c_{b},c_{w}\}$ throughout the article. Whenever unambiguous, we use the abbreviated notations $C=\{1,2,\ldots,n\}$ and $D=\{1,2,\ldots,n\}\setminus\{b,w\}$.
The Pairwise Comparison System (PCS) is the pair $(A_{b},A_{w})$, where $A_{b}=(a_{b1},a_{b2},\ldots,a_{bn})$ is the best-to-other vector and $A_{w}=(a_{1w},a_{2w},\ldots,a_{nw})^{T}$ is the other-to-worst vector. Here, $a_{ij}$ represents the relative preference of the $i^{th}$ criterion over the $j^{th}$ criterion.
Definition 1.
[32] A PCS $(A_{b},A_{w})$ is said to be consistent if $a_{bi}\times a_{iw}=a_{bw}$ for all $i\in D$.
Theorem 1.
[47] The system of linear equations
$\begin{split}&\frac{w_{b}}{w_{i}}=a_{bi},\quad\frac{w_{i}}{w_{w}}=a_{iw},\quad\frac{w_{b}}{w_{w}}=a_{bw},\ i\in D,\\ &w_{1}+w_{2}+\ldots+w_{n}=1\end{split}$ (1)
has a solution if and only if $(A_{b},A_{w})$ is consistent. Also, if solution exists, then it is unique and is given by
$w_{i}=\frac{a_{iw}}{\displaystyle\sum_{j\in C}a_{jw}}=\frac{1}{a_{bi}\displaystyle\sum_{j\in C}\frac{1}{a_{bj}}},\ i\in C.$ (2)
2.2 Nonlinear BWM
The BWM is an MCDM technique that derives criteria weights through pairwise comparisons between the best (most preferable), the worst (least preferable) and other criteria [32]. The steps of the BWM are as follows.
Step 1: Formation of the set of decision criteria $C=\{c_{1},c_{2},\ldots,c_{n}\}$.
Furthermore, We adopt the notation $D=\{c_{1},c_{2},\ldots,c_{n}\}\setminus\{c_{b},c_{w}\}$ throughout this work. When no ambiguity arises, we simplify the notation to $C=\{1,2,\ldots,n\}$ and $D=\{1,2,\ldots,n\}\setminus\{b,w\}.$
Step 2: Selection of the best criterion $c_{b}$ and the worst criterion $c_{w}$ from $C$.
Step 3: Determination of the best-to-other vector $A_{b}=(a_{b1},a_{b2},\ldots,a_{bn})$ and the other-to-worst vector $A_{w}=(a_{1w},a_{2w},\ldots,a_{nw})^{T}$.
The preferences $a_{ij}$ are typically expressed as linguistic terms, which are then quantified using established scales as shown in Table 1.
Table 1: Quantification of linguistic terms using different scales
Linguistic term Saaty Salo-Hämäläinen Lootsma Donegan-Dodd-McMaster
scale [37] scale [21] scale [21] scale (7-based) [14]
Indifference 11 11 11 11
- 22 $1.2222$ $\sqrt{2}$ $1.1257$
Moderate preference 33 $1.5$ 22 $1.2715$
- 44 $1.8571$ $2\sqrt{2}$ $1.4470$
Strong preference 55 $2.3333$ 44 $1.6684$
- 66 33 $4\sqrt{2}$ $1.9670$
Very strong preference 77 44 88 $2.4142$
- 88 $5.6667$ $8\sqrt{2}$ $3.2289$
Extreme preference 99 99 $16$ $5.8284$
Step 4: Computation of optimal weights using a nonlinear optimization model.
Consider the following minimization problem.
$\begin{split}&\min\epsilon\\ &\text{subject to:}\\ &\left|\frac{w_{b}}{w_{i}}-a_{bi}\right|\leq\epsilon,\quad\left|\frac{w_{i}}{w_{w}}-a_{iw}\right|\leq\epsilon,\quad\left|\frac{w_{b}}{w_{w}}-a_{bw}\right|\leq\epsilon,\\ &w_{1}+w_{2}+\ldots+w_{n}=1,\quad w_{j}\geq 0\text{ for all }i\in D\text{ and }j\in C.\end{split}$ (3)
Problem (3) has optimal solution(s) of the form $(w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*},\epsilon^{*})$. For each optimal solution, $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$ is an optimal weight set, while $\epsilon^{*}$ indicates its accuracy. Since $\epsilon^{*}$ is the optimal objective value, it remains the same for all optimal weight sets.
To address the non-uniqueness of optimal solutions in problem (3), Rezaei [33] employed interval-analysis, observing that the set of all optimal weights for each criterion is an interval. These optimal interval-weights can be obtained using the following optimization problems.
$\begin{split}&\min w_{k}\\ &\text{subject to:}\\ &\left|\frac{w_{b}}{w_{i}}-a_{bi}\right|\leq\epsilon^{*},\quad\left|\frac{w_{i}}{w_{w}}-a_{iw}\right|\leq\epsilon^{*},\quad\left|\frac{w_{b}}{w_{w}}-a_{bw}\right|\leq\epsilon^{*},\\ &w_{1}+w_{2}+\ldots+w_{n}=1,\quad w_{j}\geq 0\text{ for all }i\in D\text{ and }j\in C.\end{split}$ (4)
$\begin{split}&\max w_{k}\\ &\text{subject to:}\\ &\left|\frac{w_{b}}{w_{i}}-a_{bi}\right|\leq\epsilon^{*},\quad\left|\frac{w_{i}}{w_{w}}-a_{iw}\right|\leq\epsilon^{*},\quad\left|\frac{w_{b}}{w_{w}}-a_{bw}\right|\leq\epsilon^{*},\\ &w_{1}+w_{2}+\ldots+w_{n}=1,\quad w_{j}\geq 0\text{ for all }i\in D\text{ and }j\in C.\end{split}$ (5)
Problems (4) and (5) are optimization problems having $n$ variables, where the Greatest Lower Bound (GLB) and the Least Upper Bound (LUB) of the optimal interval-weight for criterion $c_{k}$ serve as the respective optimal objective values, i.e., if $w_{k}^{\prime*}$ and $w_{k}^{\prime\prime*}$ denote the optimal objective values of problems (4) and (5) respectively, then the optimal interval-weight for $c_{k}$ is $w_{k}^{\prime*},w_{k}^{\prime\prime*}]$.
The effectiveness of an MCDM method depends on the decision data that is often inconsistent because of human engagement. A key requirement for any rigorous MCDM methodology is the ability to assess and quantify these inconsistencies. In the BWM, this assessment is performed using the Consistency Ratio (CR) defined as
$\text{CR}=\frac{\epsilon^{*}}{\text{Consistency Index (CI)}},$ (6)
where CI $\sup\{\epsilon^{*}:\epsilon^{*}$ is the optimal objective value of problem (3) for some $(A_{b},A_{w})$ having the given value of $a_{BW}$} [32]. So, CI is a function of $a_{bw}$. The values of CI for the Saaty scale are given in Table 2.
Table 2: The values of CI for the Saaty scale [32]
$a_{bw}$ 22 33 44 55 66 77 88 99
CI $0.4384$ 11 $1.6277$ $2.2984$ 33 $3.7250$ $4.4688$ $5.2279$
2.3 An Analytical Framework for the Nonlinear BWM
Wu et al. [47] proposed an analytical approach to derive optimal interval weights without requiring optimization software. They also introduced a secondary objective function to select the best optimal weight set from the collection of all optimal weight sets.
2.3.1 Calculation of Interval-Weights
Consider the following optimization model, driven by the optimal modification of PCS.
$\begin{split}&\min\eta\\ &\text{subject to:}\\ &\left|\tilde{a}_{bi}-a_{bi}\right|\leq\eta,\quad\left|\tilde{a}_{iw}-a_{iw}\right|\leq\eta,\quad\left|\tilde{a}_{bw}-a_{bw}\right|\leq\eta,\\ &\tilde{a}_{bi}\times\tilde{a}_{iw}=\tilde{a}_{bw},\quad\tilde{a}_{bi},\tilde{a}_{iw},\tilde{a}_{bw}\geq 0\text{ for all }i\in D.\end{split}$ (7)
Problem (7) has optimal solution(s) of the form $(\tilde{a}_{bi}^{*},\tilde{a}_{iw}^{*},\tilde{a}_{bw}^{*},\eta^{*})$, where $i\in D$. Each optimal solution, along with $\tilde{a}_{bb}^{*}=\tilde{a}_{ww}^{*}=1$, leads to a consistent PCS $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$, referred to as the optimally modified PCS and $\eta^{*}$ indicates its the accuracy. Since $\eta^{*}$ is the optimal objective value, it remains the same for all $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$.
Wu et al. [47] established that
1. $\epsilon^{*}=\eta^{*}$ 2. for each $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$, there exists a unique $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ satisfying the relation $w_{i}^{*}=\frac{\tilde{a}_{iw}^{*}}{\sum_{i=1}^{n}\tilde{a}_{jw}^{*}}$. (8)
This implies that problem (7) is an equivalent to problem (3). Consequently, the analytical solution of problem (3) can be obtained by solving problem (7) analytically. To describe the analytical solution of Problem (7), some mathematical symbols must first be defined.
Let $D_{1}=\{i\in D:a_{bi}\times a_{iw}<a_{bw}\}$, $D_{2}=\{i\in D:a_{bi}\times a_{iw}>a_{bw}\}$ and $D_{3}=\{i\in D:a_{bi}\times a_{iw}=a_{bw}\}$.
Fix $i\in D$. Then there are three possibilities.
  1. (i)
    $i\in D_{1}$
    Consider the quadratic equation
    $(a_{bi}+x)\times(a_{iw}+x)=a_{bw}-x.$ (9)
    Let $f(x)=(a_{bi}+x)\times(a_{iw}+x)$ and $g(x)=a_{bw}-x$, where $x\in\mathbb{R}$. Note that $f(0)=a_{bi}\times a_{iw}<a_{bw}=g(0)$ and $f(a_{bw})=(a_{bi}+a_{bw})\times(a_{iw}+a_{bw})>0=g(a_{bw})$. So, by IVT, there exist $0<c<a_{bw}$ such that $f(c)=g(c)$, i.e., $c$ is a positive root of equation (9). Let $\epsilon_{i}$ be the smallest positive root of equation (9). Then
    $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$ (10)
    From the above discussion, it follows that $\epsilon_{i}<a_{bw}$.
  2. (ii)
    $i\in D_{2}$
    Consider the quadratic equation
    $(a_{bi}-x)\times(a_{iw}-x)=a_{bw}+x.$ (11)
    Let $f(x)=(a_{bi}-x)\times(a_{iw}-x)$ and $g(x)=a_{bw}+x$, where $x\in\mathbb{R}$. Let $a=\min\{a_{bi},a_{iw}\}$ Note that $f(0)=a_{bi}\times a_{iw}>a_{bw}=g(0)$ and $f(a)=0<a_{bw}+a=g(a)$. So, by IVT, there exist $0<c<a$ such that $f(c)=g(c)$, i.e., $c$ is a positive root of equation (11). Let $\epsilon_{i}$ be the smallest positive root of equation (11). Then
    $(a_{bi}-\epsilon_{i})\times(a_{iw}-\epsilon_{i})=a_{bw}+\epsilon_{i}.$ (12)
    From the above discussion, it follows that $\epsilon_{i}<a$, i.e., $\epsilon_{i}<a_{bi}$ and $\epsilon_{i}<a_{iw}$.
  3. (iii)
    $i\in D_{3}$
    In this case, take $\epsilon_{i}=0$.
So, in any case, we get
$\epsilon_{i}=\bigg{|}\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg{|}.$ (13)
Now, fix $i,j\in D$. Then there are three possibilities.
  1. (i)
    $a_{bi}\times a_{iw}<a_{bj}\times a_{jw}$
    In this case, take $\epsilon_{i,j}=\frac{a_{bj}\times a_{jw}-a_{bi}\times a_{iw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}.$ This gives
    $(a_{bi}+\epsilon_{i,j})\times(a_{iw}+\epsilon_{i,j})=(a_{bj}-\epsilon_{i,j})\times(a_{jw}-\epsilon_{i,j}).$ (14)
    Note that $\epsilon_{i,j}<a_{bj}$ and $\epsilon_{i,j}<a_{jw}$.
  2. (ii)
    $a_{bi}\times a_{iw}>a_{bj}\times a_{jw}$
    In this case, take $\epsilon_{i,j}=\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}.$ This gives
    $(a_{bi}-\epsilon_{i,j})\times(a_{iw}-\epsilon_{i,j})=(a_{bj}+\epsilon_{i,j})\times(a_{jw}+\epsilon_{i,j}).$ (15)
    Note that $\epsilon_{i,j}<a_{bi}$ and $\epsilon_{i,j}<a_{iw}$.
  3. (iii)
    $a_{bi}\times a_{iw}=a_{bj}\times a_{jw}$
    In this case, take $\epsilon_{i,j}=0$.
So, in any case, we get
$\epsilon_{i,j}=\bigg{|}\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}\bigg{|}.$ (16)
Let $i_{1}\in D_{1}$ and $j_{0}\in D_{2}$ be such that $\epsilon_{i_{1}}=\max\{\epsilon_{i}:i\in D_{1}\}$ and $\epsilon_{i_{2}}=\max\{\epsilon_{i}:i\in D_{2}\}$ respectively. Then, by [47, Proposition 3], we get
$\epsilon^{*}=\begin{cases}\epsilon_{i_{1}}\quad\ \ \ \text{if }(a_{bi_{2}}-\epsilon_{i_{1}})\times(a_{i_{2}w}-\epsilon_{i_{1}})\leq a_{bw}-\epsilon_{i_{1}},\\ \epsilon_{i_{2}}\quad\ \ \ \text{if }(a_{bi_{1}}+\epsilon_{i_{2}})\times(a_{i_{1}w}+\epsilon_{i_{2}})\geq a_{bw}+\epsilon_{i_{2}},\\ \epsilon_{i_{1},i_{2}}\quad\text{otherwise}.\end{cases}$ (17)
Now, by [47, Theorem 3], the collection of all optimally modified PCS is
$\tilde{a}_{bw}^{*}=\begin{cases}a_{bw}-\epsilon^{*}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if }\epsilon^{*}=\epsilon_{i_{1}},\\ a_{bw}+\epsilon^{*}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if }\epsilon^{*}=\epsilon_{i_{2}},\\ (a_{bi_{2}}-\epsilon_{i_{1},i_{2}})\times(a_{i_{2}w}-\epsilon_{i_{1},i_{2}})\quad\text{if }\epsilon^{*}=\epsilon_{i_{1},i_{2}},\end{cases}$ (18a)
$\tilde{a}_{iw}^{*}\in\bigg{[}\max\biggl{\{}a_{iw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}+\epsilon^{*}}\biggr{\}},\min\biggl{\{}a_{iw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}-\epsilon^{*}}\biggr{\}}\bigg{]}\text{ with }\tilde{a}_{bi}^{*}=\frac{\tilde{a}_{bw}^{*}}{\tilde{a}_{iw}^{*}},\text{ where }i\in D.$ (18b)
Now, from [47, Theorem 4], the collection of all optimal weights of criterion $c_{i}$ is ${w_{i}^{l}}^{*},{w_{i}^{u}}^{*}]$, where
${w_{i}^{l}}^{*}=\frac{\inf\{\tilde{a}_{iw}^{*}\}}{\inf\{\tilde{a}_{iw}^{*}\}+\displaystyle\sum_{\begin{subarray}{c}j\in C\\ j\neq i\end{subarray}}\sup\{\tilde{a}_{jw}^{*}\}},\quad{w_{i}^{u}}^{*}=\frac{\sup\{\tilde{a}_{iw}^{*}\}}{\sup\{\tilde{a}_{iw}^{*}\}+\displaystyle\sum_{\begin{subarray}{c}j\in C\\ j\neq i\end{subarray}}\inf\{\tilde{a}_{jw}^{*}\}},$ (19)
where $i\in C$, and thus,
$\begin{split}{w_{b}^{l}}^{*}&=\frac{\tilde{a}_{bw}^{*}}{1+\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D}\min\biggl{\{}a_{jw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}-\epsilon^{*}}\biggr{\}}},\quad{w_{b}^{u}}^{*}=\frac{\tilde{a}_{bw}^{*}}{1+\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D}\max\biggl{\{}a_{jw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}+\epsilon^{*}}\biggr{\}}},\\ {w_{w}^{l}}^{*}&=\frac{1}{1+\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D}\min\biggl{\{}a_{jw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}-\epsilon^{*}}\biggr{\}}},\quad{w_{w}^{u}}^{*}=\frac{1}{1+\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D}\max\biggl{\{}a_{jw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}+\epsilon^{*}}\biggr{\}}},\\ {w_{i}^{l}}^{*}&=\frac{\max\biggl{\{}a_{iw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}+\epsilon^{*}}\biggr{\}}}{1+\tilde{a}_{bw}^{*}+\max\biggl{\{}a_{iw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}+\epsilon^{*}}\biggr{\}}+\displaystyle\sum_{\begin{subarray}{c}j\in D\\ j\neq i\end{subarray}}\min\biggl{\{}a_{jw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}-\epsilon^{*}}\biggr{\}}},\\ {w_{i}^{u}}^{*}&=\frac{\min\biggl{\{}a_{iw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}-\epsilon^{*}}\biggr{\}}}{1+\tilde{a}_{bw}^{*}+\min\biggl{\{}a_{iw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}-\epsilon^{*}}\biggr{\}}+\displaystyle\sum_{\begin{subarray}{c}j\in D\\ j\neq i\end{subarray}}\max\biggl{\{}a_{jw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}+\epsilon^{*}}\biggr{\}}},\quad\text{where }i\in D.\end{split}$ (20)
2.3.2 A Secondary Objective Function to Obtain a Unique Weight Set
Since the nonlinear BWM may yield multiple optimal weight sets—and while such multiplicity is desirable in some cases—DMs usually prefer a unique weight set. To address this, Wu et al. [47] introduced a secondary objective function to obtain the best optimal weight set.
In this approach, an optimally modified PCS having the minimum value of $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}$ for all $i\in C$ is first selected as the best optimally modified PCS. As shown in [47, Theorem 5], for $i\in D$, the minimum possible value of $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}$ is the optimal objective value of minimization problem
$\begin{split}&\min\eta_{i}\\ &\text{subject to:}\\ &\tilde{a}_{bi}-a_{bi}=\eta_{bi},\quad\tilde{a}_{iw}-a_{iw}=\eta_{iw},\quad(a_{bi}+\eta_{bi})\times(a_{iw}+\eta_{iw})=\tilde{a}_{bw}^{*},\\ &0\leq\eta_{bi},\eta_{iw}\leq\eta_{i}\end{split}$ (21)
if $a_{bi}\times a_{iw}\leq\tilde{a}_{bw}^{*}$ and the optimal objective value of minimization problem
$\begin{split}&\min\eta_{i}\\ &\text{subject to:}\\ &a_{bi}-\tilde{a}_{bi}=\eta_{bi},\quad a_{iw}-\tilde{a}_{iw}=\eta_{iw},\quad(a_{bi}-\eta_{bi})\times(a_{iw}-\eta_{iw})=\tilde{a}_{bw}^{*},\\ &0\leq\eta_{bi},\eta_{iw}\leq\eta_{i}\end{split}$ (22)
if $a_{bi}\times a_{iw}>\tilde{a}_{bw}^{*}$. Furthermore, [47, Theorem 6] states that the PCS
[…]
where $\tilde{a}_{bw}^{*}$ is as defined in equation (18) and
$\eta_{i}^{*}=\bigg{|}\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4(a_{bi}\times a_{iw}-\tilde{a}_{bw}^{*})}}{2}\bigg{|},\quad i\in D,$ (24)
is the only optimally modified PCS. Now, from this PCS, the best optimal weight set can be obtained using Theorem 1.
3 Research Gap
Although the analytical framework proposed by Wu et al. [47] significantly improves the model’s efficiency, some research gaps remain unaddressed.
(i) Incompatibility of the framework with some scales:
The framework is fully compatible with the Saaty scale but proves incompatible with some alternative scales, including the Salo-Hämäläinen scale, the Lootsma scale and the Donegan-Dodd-McMaster scale.
Example 1: Let $C=\{c_{1},c_{2},\ldots,c_{5}\}$ be the set of decision criteria with $c_{1}$ as the best and $c_{5}$ as the worst criterion. Let $A_{b}=(1,9,3,1.8571,9)$ be the best-to-other vector and $A_{w}=(9,1.5,4,3,1)^{T}$ be the other-to-worst vector, obtained by quantifying the preferences given in the form of linguistic terms using the Salo-Hämäläinen scale.
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=0.4056$, $\epsilon_{3}=0.3944$ and $\epsilon_{4}=0.5363$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=0.5163$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})<a_{15}-\epsilon_{4}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4}=0.5363$. The optimal interval-weights computed using equation (20) are given in Table 3.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=8.4638$. By equation (24), we get $\eta_{2}^{*}=0.5038$, $\eta_{3}^{*}=0.5481$ and $\eta_{4}^{*}=0.5363$. From equation (23), it follows that the best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,8.4962,2.4519,2.3934,8.4638),\quad\tilde{A}_{w}^{*}=(8.4638,0.9962,3.4519,3.5363,1)^{T}.$
The best optimal weight set calculated using Theorem 1 is presented in Table 3.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 3. The best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,8.4999,2.4578,2.3993,8.4987),\quad\tilde{A}_{w}^{*}=(8.4987,0.9999,3.4578,3.5422,1)^{T}.$
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 3.
Table 3: Weights and $\epsilon^{*}$ for Example 1
Analytical approach [47] Actual value
Criterion Interval-weight Best weight Interval-weight Best weight
$c_{1}$ $[0.4854,0.4857]$ $0.4851$ $[0.4855,0.4868]$ $0.4857$
$c_{2}$ $[0.0554,0.0573]$ $0.0571$ $[0.0549,0.0574]$ $0.0571$
$c_{3}$ $[0.1983,0.1974]$ $0.1978$ $[0.1975,0.1981]$ $0.1976$
$c_{4}$ $[0.2028,0.2029]$ $0.2027$ $[0.2024,0.2029]$ $0.2024$
$c_{5}$ $[0.0573,0.0574]$ $0.0573$ $[0.0571,0.0573]$ $0.0572$
$\epsilon^{*}$ $0.5363$ $0.5422$
Example 2: Let $C=\{c_{1},c_{2},\ldots,c_{5}\}$ be the set of decision criteria with $c_{1}$ as the best and $c_{5}$ as the worst criterion. Let $A_{b}=(1,16,4\sqrt{2},2\sqrt{2},16)$ be the best-to-other vector and $A_{w}=(16,2\sqrt{2},4\sqrt{2},\sqrt{2},1)^{T}$ be the other-to-worst vector, obtained by quantifying the preferences given in the form of linguistic terms using the Lootsma scale.
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=1.6054$, $\epsilon_{3}=1.4764$ and $\epsilon_{4}=1.7228$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=1.7882$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{2})\times(a_{45}+\epsilon_{2})<a_{15}+\epsilon_{2}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,2}=1.7882$. The optimal interval-weights computed using equation (20) are given in Table 4.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=14.7841$. By equation (24), we get $\eta_{2}^{*}=1.7882$, $\eta_{3}^{*}=1.8118$ and $\eta_{4}^{*}=1.7882$. From equation (23), it follows that the best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,14.2118,3.8451,4.6166,14.7841),\quad\tilde{A}_{w}^{*}=(14.7841,1.0403,3.8451,3.2024,1)^{T}.$
The best optimal weight set calculated using Theorem 1 is presented in Table 4.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 4. The best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,14.2179,3.8569,4.6283,14.8760),\quad\tilde{A}_{w}^{*}=(14.8760,1.0463,3.8569,3.2141,1)^{T}.$
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 4.
Table 4: Weights and $\epsilon^{*}$ for Example 2
Analytical approach [47] Actual value
Criterion Interval-weight Best weight Interval-weight Best weight
$c_{1}$ $[0.6199,0.6187]$ $0.6193$ $[0.6200,0.6205]$ $0.6200$
$c_{2}$ $[0.0436,0.0435]$ $0.0436$ $[0.0429,0.0437]$ $0.0436$
$c_{3}$ $[0.1619,0.1602]$ $0.1611$ $[0.1607,0.1609]$ $0.1607$
$c_{4}$ $[0.1343,0.1340]$ $0.1341$ $[0.1340,0.1341]$ $0.1340$
$c_{5}$ $[0.0419,0.0418]$ $0.0419$ $[0.0416,0.0417]$ $0.0417$
$\epsilon^{*}$ $1.7882$ $1.7999$
Example 3: Let $C=\{c_{1},c_{2},\ldots,c_{5}\}$ be the set of decision criteria with $c_{1}$ as the best and $c_{5}$ as the worst criterion. Let $A_{b}=(1,5.8284,3.2289,1,5.8284)$ be the best-to-other vector and $A_{w}=(5.8284,1.967,3.2289,1.967,1)^{T}$ be the other-to-worst vector, obtained by quantifying the preferences given in the form of linguistic terms using the Donegan-Dodd-McMaster scale.
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=0.6959$, $\epsilon_{3}=0.6781$ and $\epsilon_{4}=0.8086$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=0.8825$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{2})\times(a_{45}+\epsilon_{2})<a_{15}+\epsilon_{2}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,2}=0.8825$. The optimal interval-weights computed using equation (20) are given in Table 5.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=5.3640$. By equation (24), we get $\eta_{2}^{*}=0.8825$, $\eta_{3}^{*}=0.9129$ and $\eta_{4}^{*}=0.8825$. From equation (23), it follows that the best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,4.9459,2.316,1.8825,5.3640),\quad\tilde{A}_{w}^{*}=(5.3640,1.0845,2.316,2.8495,1)^{T}.$
The best optimal weight set calculated using Theorem 1 is presented in Table 5.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 5. The best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,4.9577,2.3314,1.8975,5.4354),\quad\tilde{A}_{w}^{*}=(5.4354,1.0963,2.3314,2.8645,1)^{T}.$
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 5.
Table 5: Weights and $\epsilon^{*}$ for Example 3
Analytical approach [47] Actual value
Criterion Interval-weight Best weight Interval-weight Best weight
$c_{1}$ $[0.4263,0.4242]$ $0.4252$ $[0.4269,0.4280]$ $0.4270$
$c_{2}$ $[0.0862,0.0858]$ $0.0860$ $[0.0842,0.0866]$ $0.0861$
$c_{3}$ $[0.1856,0.1817]$ $0.1836$ $[0.1831,0.1836]$ $0.1832$
$c_{4}$ $[0.2264,0.2254]$ $0.2259$ $[0.2250,0.2255]$ $0.2251$
$c_{5}$ $[0.0795,0.0791]$ $0.0793$ $[0.0785,0.0787]$ $0.0786$
$\epsilon^{*}$ $0.8825$ $0.8975$
(ii) Incompatibility of the framework with the Saaty scale in the presence of multiple DMs:
The framework does not work even with the Saaty scale in group decision contexts where the Aggregation of Individual Judgements (AIJ) approach is employed [6].
Example 4: Let $C=\{c_{1},c_{2},\ldots,c_{5}\}$ be the set of decision criteria with $c_{1}$ as the best and $c_{5}$ as the worst criterion. Let $E_{1}$ and $E_{2}$ be two homogeneous DMs. Let $(A_{b})_{1}=(1,5,1,3,7)$ be the best-to-other vector and $(A_{w})_{1}=(7,2,7,1,1)^{T}$ be the other-to-worst vector for $E_{1}$, and let $(A_{b})_{2}=(1,2,5,2,7)$ be the best-to-other vector and $(A_{w})_{2}=(7,5,3,3,1)^{T}$ be the other-to-worst vector for $E_{2}$. These vectors are obtained by quantifying the preferences given in the form of linguistic terms using the Saaty scale.
Now, we aggregate these individual judgments using the geometric mean method. Thus, we get $A_{b}=(1,\sqrt{10},\sqrt{5},\sqrt{6},7)$ as the aggregated best-to-other vector and $A_{w}=(7,\sqrt{10},\sqrt{21},\sqrt{3},1)^{T}$ as the aggregated other-to-worst vector.
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=0.4355$, $\epsilon_{3}=0.4401$ and $\epsilon_{4}=0.4865$. This gives $i_{1}=4$ and $i_{2}=3$. Now, by equation (16), we get $\epsilon_{4,3}=0.5458$. Note that $(a_{13}-\epsilon_{4})\times(a_{35}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{3})\times(a_{45}+\epsilon_{3})<a_{15}+\epsilon_{3}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,3}=0.5458$. The optimal interval-weights computed using equation (20) are given in Table 6.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=6.8230$. By equation (24), we get $\eta_{2}^{*}=0.5502$, $\eta_{3}^{*}=0.5458$ and $\eta_{4}^{*}=0.5458$. From equation (23), it follows that the best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,2.6121,1.6902,2.9953,6.8230),\quad\tilde{A}_{w}^{*}=(6.8230,2.6121,4.0367,2.2779,1)^{T}.$
The best optimal weight set calculated using Theorem 1 is presented in Table 6.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{2}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 6. The best optimally modified PCS is
$\tilde{A}_{b}^{*}=(1,2.6143,1.6923,2.9975,6.8344),\quad\tilde{A}_{w}^{*}=(6.8344,2.6143,4.0388,2.2801,1)^{T}.$
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 6.
Table 6: Weights and $\epsilon^{*}$ for Example 4
Analytical approach [47] Actual value
Criterion Interval-weight Best weight Interval-weight Best weight
$c_{1}$ $[0.4099,0.4047]$ $0.4074$ $[0.4074,0.4077]$ $0.4076$
$c_{2}$ $[0.1615,0.1506]$ $0.1559$ $[0.1558,0.1560]$ $0.1559$
$c_{3}$ $[0.2425,0.2394]$ $0.2410$ $[0.2407,0.2413]$ $0.2409$
$c_{4}$ $[0.1369,0.1351]$ $0.1360$ $[0.1359,0.1360]$ $0.1360$
$c_{5}$ $[0.0601,0.0593]$ $0.0597$ $[0.0596,0.0597]$ $0.0596$
$\epsilon^{*}$ $0.5458$ $0.5480$
(iii) Non-well-defined values of CI for the Saaty scale:
The CI for the Saaty scale, presented in Table 2, does not serve as an upper bound for the set of $\epsilon^{*}$ corresponding to the PCSs having the given $a_{bw}$, and thus, is not well-defined.
Example 5: Let $\{c_{1},c_{2},c_{3},c_{4}\}$ be the set of decision criteria with $c_{1}$ as the best and $c_{4}$ as the worst criterion. Let $A_{b}=(1,1,2,2)$ be the best-to-other vector and $A_{w}=(2,1,2,1)^{T}$ be the other-to-worst vector, obtained by quantifying the preferences given in the form of linguistic terms using the Saaty scale.
Here, we get a unique optimal weight set, which along with $\epsilon^{*}$, is given in Table 7.
Note that $\epsilon^{*}=0.5>0.4384=\text{CI}_{a_{bw}=2}$. This implies that the values of CI for the Saaty scale given in Table 2 are not well-defined.
Table 7: Weights and $\epsilon^{*}$ for Example 5
Criterion Weight
$c_{1}$ $0.36$
$c_{2}$ $0.24$
$c_{3}$ $0.24$
$c_{4}$ $0.16$
$\epsilon^{*}$ $0.5$
(iv) Weight differences among criteria with equal preference in the presence of multiple best/worst criteria:
In instances where multiple criteria are equally qualified as the best (or worst), the conventional approach involves arbitrarily selecting one as the best (or worst). However, the chosen one might get a different weight than the others even though they are of equal preference.
Example 6: Let $\{c_{1},c_{2},c_{3},c_{4}\}$ be the set of decision criteria with $c_{1}$ and $c_{2}$ as the best criteria and $c_{4}$ as the worst criterion. Here, we select $c_{1}$ as the best criterion and proceed accordingly. Let $A_{b}=(1,1,2,7)$ be the best-to-other vector and $A_{w}=(7,7,3,1)^{T}$ be the other-to-worst vector, obtained by quantifying the preferences given in the form of linguistic terms using the Saaty scale.
The interval-weights, the best weight set and $\epsilon^{*}$ are given in Table 8.
Note that for $c_{1}$ and $c_{2}$, neither the interval-weights nor the optimal weights coincide, despite both criteria being of equal preferences.
Table 8: Weights and $\epsilon^{*}$ for Example 6
Criterion Interval-weight Best weight
$c_{1}$ $[0.3765,0.3833]$ $0.3803$
$c_{2}$ $[0.3833,0.3944]$ $0.3882$
$c_{3}$ $[0.1741,0.1773]$ $0.1759$
$c_{4}$ $[0.0551,0.0561]$ $0.0556$
$\epsilon^{*}$ $0.1623$
4 A Generalized Analytical Framework for the Nonlinear Best-Worst Method
In this section, we propose a generalized analytical framework that is compatible with any scale and any number of DMs. We derive a formula for CI and an analytical expression for CR. Furthermore, we modify the original optimization model to ensure equal weights for criteria with identical preferences when multiple best/worst criteria are present.
4.1 Calculation of Weights
This subsection consists of two parts: weight calculation when a unique best and worst criterion exists, and weight calculation for cases with multiple best or worst criteria.
4.1.1 Calculation of Weights in the Presence of Unique Best and Worst Criterion
Proposition 1.
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then $\epsilon_{i}\leq\eta^{*}$ and $\epsilon_{i,j}\leq\eta^{*}$ for all $i,j\in D$.
Proof.
Let $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ be an optimally modified PCS. Let $|\tilde{a}_{bi}^{*}-a_{bi}|=\eta_{bi}$, $|\tilde{a}_{iw}^{*}-a_{iw}|=\eta_{iw}$ and $|\tilde{a}_{bw}^{*}-a_{bw}|=\eta_{bw}$. Therefore, $0\leq\eta_{bi},\eta_{iw},\eta_{bw}\leq\eta^{*}$. Also, $\tilde{a}_{bi}^{*}\in\{a_{bi}+\eta_{bi},a_{bi}-\eta_{bi}\}$, $\tilde{a}_{iw}^{*}\in\{a_{iw}+\eta_{iw},a_{iw}-\eta_{iw}\}$ and $\tilde{a}_{bw}^{*}\in\{a_{bw}+\eta_{bw},a_{bw}-\eta_{bw}\}$.
Fix $i\in D$. If $i\in D_{3}$, then by equation (13), $\epsilon_{i}=0$ and we are done. If $i\in D_{1}$, then by equation (10), we have $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$. Now, to prove $\epsilon_{i}\leq\eta^{*}$, it suffices to show that at least one of the inequalities $\epsilon_{i}\leq\eta_{bi}$, $\epsilon_{i}\leq\eta_{iw}$ or $\epsilon_{i}\leq\eta_{bw}$ holds. Suppose, if possible, neither of these inequalities hold. Then we get $a_{bi}+\epsilon_{i}>\tilde{a}_{bi}^{*}$, $a_{iw}+\epsilon_{i}>\tilde{a}_{iw}^{*}$ and $a_{bw}-\epsilon_{i}<\tilde{a}_{bw}^{*}$. This implies $\tilde{a}_{bi}^{*}\times\tilde{a}_{iw}^{*}<\tilde{a}_{bw}^{*}$, which is contradiction as $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent. Thus, $\epsilon_{i}\leq\eta^{*}$. If $i\in D_{2}$, then the result follows by applying the same argument as above.
Fix $i,j\in D$. If $a_{bi}\times a_{iw}=a_{bj}\times a_{jw}$. then by equation (16), $\epsilon_{i,j}=0$ and we are done. If $a_{bi}\times a_{iw}<a_{bj}\times a_{jw}$, then by equation (14), we have $(a_{bi}+\epsilon_{i,j})\times(a_{iw}+\epsilon_{i,j})=(a_{bj}-\epsilon_{i,j})\times(a_{jw}-\epsilon_{i,j})$. Now, to prove $\epsilon_{i,j}\leq\eta^{*}$, it suffices to show that at least one of the inequalities $\epsilon_{i,j}\leq\eta_{bi}$, $\epsilon_{i,j}\leq\eta_{iw}$, $\epsilon_{i,j}\leq\eta_{bj}$ or $\epsilon_{i,j}\leq\eta_{jw}$ holds. Suppose, if possible, neither of these inequalities hold. Then we get $a_{bi}+\epsilon_{i,j}>\tilde{a}_{bi}^{*}$, $a_{iw}+\epsilon_{i,j}>\tilde{a}_{iw}^{*}$, $a_{bj}-\epsilon_{i,j}<\tilde{a}_{bj}^{*}$ and $a_{jw}-\epsilon_{i,j}<\tilde{a}_{jw}^{*}$. This implies $\tilde{a}_{bi}^{*}\times\tilde{a}_{iw}^{*}<\tilde{a}_{bj}^{*}\times\tilde{a}_{jw}^{*}$, which is contradiction as $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent. Thus, $\epsilon_{i,j}\leq\eta^{*}$. Hence the proof. ∎
Theorem 2.
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then the following statements hold.
  1. 1.
    If $\epsilon_{i_{0}}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ for some $i_{0}\in D_{1}$, then $\eta^{*}=\epsilon_{i_{0}}$. Also, $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
    $\begin{split}&\tilde{a}_{bi}=\frac{a_{bi}-a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2},\\ &\tilde{a}_{iw}=\frac{-a_{bi}+a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\quad\text{for all }i\in D,\\ &\tilde{a}_{bw}=a_{bw}-\epsilon_{i_{0}}\end{split}$ (25)
    is an optimally modified PCS.
  2. 2.
    If $\epsilon_{j_{0}}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ for some $j_{0}\in D_{2}$, then $\eta^{*}=\epsilon_{j_{0}}$. Also, $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
    $\begin{split}&\tilde{a}_{bi}=\frac{a_{bi}-a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}+\epsilon_{j_{0}})}}{2},\\ &\tilde{a}_{iw}=\frac{-a_{bi}+a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}+\epsilon_{j_{0}})}}{2}\quad\text{for all }i\in D,\\ &\tilde{a}_{bw}=a_{bw}+\epsilon_{j_{0}}\end{split}$ (26)
    is an optimally modified PCS.
  3. 3.
    If $\epsilon_{i_{0},j_{0}}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ for some $i_{0},j_{0}\in D$, then $\eta^{*}=\epsilon_{i_{0},j_{0}}$. Also, $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
    $\begin{split}&\tilde{a}_{bi}=\frac{a_{bi}-a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2},\\ &\tilde{a}_{iw}=\frac{-a_{bi}+a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{for all }i\in D,\\ &\tilde{a}_{bw}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})\end{split}$ (27)
    is an optimally modified PCS.
Proof.
First, assume that $\epsilon_{i_{0}}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ for some $i_{0}\in D_{1}$. It is easy to verify that $(\tilde{A}_{b},\tilde{A}_{w})$ given in equation (25) is consistent. Now, if we prove that
$|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}},\quad|\tilde{a}_{iw}-a_{iw}|\leq\epsilon_{i_{0}},\quad|\tilde{a}_{bw}-a_{bw}|\leq\epsilon_{i_{0}}$
for all $i\in D$, then it will imply that $\eta^{*}\leq\epsilon_{i_{0}}$. This, along with Proposition 1, gives $\eta^{*}=\epsilon_{i_{0}}$, and consequently, $(\tilde{A}_{b},\tilde{A}_{w})$ is an optimally modified PCS.
Now, $|\tilde{a}_{bw}-a_{bw}|=|-\epsilon_{i_{0}}|=\epsilon_{i_{0}}$. Also,
$|\tilde{a}_{bi}-a_{bi}|=|\tilde{a}_{iw}-a_{iw}|=\bigg{|}\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{|}.$
Fix $i\in D$. Then there are two possibilities.
  1. (i)
    $a_{bi}\times a_{iw}\leq a_{bw}-\epsilon_{i_{0}}$
    Here, we get $(a_{bi}+a_{iw})\leq\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}$. Therefore,
    $\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\geq 0,$
    and thus,
    $|\tilde{a}_{bi}-a_{bi}|=|\tilde{a}_{iw}-a_{iw}|=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}.$ (28)
    Note that
    $\begin{split}\bigg{(}a_{bi}&+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}\\ &\times\bigg{(}a_{iw}+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}=a_{bw}-\epsilon_{i_{0}}.\end{split}$ (29)
    Since $a_{bi}\times a_{iw}\leq a_{bw}-\epsilon_{i_{0}}<a_{bw}$, by equation (10), we have $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$. Now, $\epsilon_{i}\leq\epsilon_{i_{0}}$ implies that $a_{bw}-\epsilon_{i_{0}}\leq a_{bw}-\epsilon_{i}$, i.e.,
    $\begin{split}\bigg{(}a_{bi}&+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}\\ &\times\bigg{(}a_{iw}+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\leq(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i}).\end{split}$
    This gives
    $|\tilde{a}_{bi}-a_{bi}|=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\leq\epsilon_{i}\leq\epsilon_{i_{0}}.$
    Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}}$.
  2. (ii)
    $a_{bi}\times a_{iw}>a_{bw}-\epsilon_{i_{0}}$
    Here, we get $\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}<a_{bi}+a_{iw}$. Therefore,
    $\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\geq 0,$
    and thus,
    $|\tilde{a}_{bi}-a_{bi}|=|\tilde{a}_{iw}-a_{iw}|=\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}.$ (30)
    Note that
    $\begin{split}\bigg{(}a_{bi}&-\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}\\ &\times\bigg{(}a_{iw}-\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}=a_{bw}-\epsilon_{i_{0}}.\end{split}$
    By equation (10), we have $(a_{bi_{0}}+\epsilon_{i_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0}})=a_{bw}-\epsilon_{i_{0}}$. Since $a_{bi}\times a_{iw}>a_{bw}-\epsilon_{i_{0}}$, we get $a_{bi_{0}}\times a_{i_{0}w}<a_{bi}\times a_{iw}$. So, by equation (14), we have $(a_{bi_{0}}+\epsilon_{i,i_{0}})\times(a_{i_{0}w}+\epsilon_{i,i_{0}})=(a_{bi}-\epsilon_{i,i_{0}})\times(a_{iw}-\epsilon_{i,i_{0}})$. Now, $\epsilon_{i,i_{0}}\leq\epsilon_{i_{0}}$ implies that $(a_{bi}-\epsilon_{i,i_{0}})\times(a_{iw}-\epsilon_{i,i_{0}})\leq a_{bw}-\epsilon_{i_{0}}$, i.e.,
    $\begin{split}&(a_{bi}-\epsilon_{i,i_{0}})\times(a_{iw}-\epsilon_{i,i_{0}})\\ &\quad\quad\quad\quad\quad\leq\bigg{(}a_{bi}-\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\times\bigg{(}a_{iw}-\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\bigg{)}.\end{split}$
    This gives
    $|\tilde{a}_{bi}-a_{bi}|=\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\leq\epsilon_{i,i_{0}}\leq\epsilon_{i_{0}}.$
    Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}}$.
The result can be proven using a similar argument as above if $\epsilon_{j_{0}}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ for some $j_{0}\in D_{2}$.
Now, assume that $\epsilon_{i_{0},j_{0}}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ for some $i,j\in D$. It is easy to check that $(\tilde{A}_{b},\tilde{A}_{w})$ given in equation (25) is consistent. Now, it is sufficient to prove that
$|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0},j_{0}},\quad|\tilde{a}_{iw}-a_{iw}|\leq\epsilon_{i_{0},j_{0}},\quad|\tilde{a}_{bw}-a_{bw}|\leq\epsilon_{i_{0},j_{0}}$
for all $i\in D$.
Let $\tilde{a}_{bw}-a_{bw}=\zeta$. Then there are two possibilities.
  1. (i)
    $\zeta\leq 0$
    Here, we get $\tilde{a}_{bw}\leq a_{bw}$. Therefore, $a_{bi_{0}}\times a_{i_{0}w}<a_{bw}$. So, by equation (10), we have $(a_{bi_{0}}+\epsilon_{i_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0}})=a_{bw}-\epsilon_{i_{0}}$. Also, $(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})=\tilde{a}_{bw}=a_{bw}+\zeta$. Now, $\epsilon_{i_{0}}\leq\epsilon_{i_{0},j_{0}}$ implies that $(a_{bi_{0}}+\epsilon_{i_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0}})\leq(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$, i.e., $a_{bw}-\epsilon_{i_{0}}\leq a_{bw}+\zeta$. This gives $|\tilde{a}_{bw}-a_{bw}|=-\zeta\leq\epsilon_{i_{0}}\leq\epsilon_{i_{0},j_{0}}$.
  2. (ii)
    $\zeta>0$
    Here, we get $a_{bw}<\tilde{a}_{bw}$. From equation (14), it follows that $\tilde{a}_{bw}=a_{bw}+\zeta=(a_{bj_{0}}-\epsilon_{i_{0},j_{0}})\times(a_{j_{0}w}-\epsilon_{i_{0},j_{0}})$. Therefore, $a_{bj_{0}}\times a_{j_{0}w}>a_{bw}$. So, by equation (12), we have $(a_{bj_{0}}-\epsilon_{j_{0}})\times(a_{j_{0}w}-\epsilon_{j_{0}})=a_{bw}+\epsilon_{j_{0}}$. Now, $\epsilon_{j_{0}}\leq\epsilon_{i_{0},j_{0}}$ implies that $(a_{bj_{0}}-\epsilon_{i_{0},j_{0}})\times(a_{j_{0}w}-\epsilon_{i_{0},j_{0}})\leq(a_{bj_{0}}-\epsilon_{j_{0}})\times(a_{j_{0}w}-\epsilon_{j_{0}})$, i.e., $a_{bw}+\zeta\leq a_{bw}+\epsilon_{j_{0}}$. This gives $|\tilde{a}_{bw}-a_{bw}|=\zeta\leq\epsilon_{j_{0}}\leq\epsilon_{i_{0},j_{0}}$.
Now,
$\begin{split}|\tilde{a}_{bi}-a_{bi}|&=|\tilde{a}_{iw}-a_{iw}|\\ &=\bigg{|}\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\bigg{|}.\end{split}$
Fix $i\in D$. Then there are two possibilities.
  1. (i)
    $a_{bi}\times a_{iw}\leq(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$
    Here, we get $(a_{bi}+a_{iw})\leq\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}$. So,
    $\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\geq 0,$
    and thus,
    $\begin{split}|\tilde{a}_{bi}-a_{bi}|&=|\tilde{a}_{iw}-a_{iw}|\\ &=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}.\end{split}$ (31)
    Note that
    $\begin{split}&\bigg{(}a_{bi}+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\bigg{)}\\ &\times\bigg{(}a_{iw}+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\bigg{)}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}}).\end{split}$ (32)
    Since $a_{bi}\times a_{iw}\leq(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$ and by equation (14), we have $(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})=(a_{bj_{0}}-\epsilon_{i_{0},j_{0}})\times(a_{j_{0}w}-\epsilon_{i_{0},j_{0}})$, we get $a_{bi}\times a_{iw}<a_{bj_{0}}\times a_{j_{0}w}$. So, from equation (14), we have $(a_{bi}+\epsilon_{i,j_{0}})\times(a_{iw}+\epsilon_{i,j_{0}})=(a_{bj_{0}}-\epsilon_{i,j_{0}})\times(a_{j_{0}w}-\epsilon_{i,j_{0}})$. Now, $\epsilon_{i,j_{0}}\leq\epsilon_{i_{0},j_{0}}$ implies that $(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})\leq(a_{bi}+\epsilon_{i,j_{0}})\times(a_{iw}+\epsilon_{i,j_{0}})$, i.e.,
    $\begin{split}&\bigg{(}a_{bi}+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\bigg{)}\\ &\times\bigg{(}a_{iw}+\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\bigg{)}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\leq(a_{bi}+\epsilon_{i,j_{0}})\times(a_{iw}+\epsilon_{i,j_{0}}).\end{split}$
    This gives
    $\begin{split}|\tilde{a}_{bi}-a_{bi}|&=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\\ &\leq\epsilon_{i,j_{0}}\leq\epsilon_{i_{0},j_{0}}.\end{split}$
    Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0},j_{0}}$.
  2. (ii)
    $a_{bi}\times a_{iw}>(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$
    In this case, the result can be proven by using a similar argument as in possibility (i).
Hence the proof. ∎
From equation 2, it follows that
$\begin{split}\epsilon^{*}=\eta^{*}&=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}\\ &=\max\left\{\bigg{|}\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg{|},\right.\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\bigg{|}\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}\bigg{|}:i,j\in D\right\},\end{split}$ (33)
which is analytical expression of $\epsilon^{*}$.
Proposition 2.
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, let $\eta^{*}$ be the optimal objective value of problem (7), and let $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ be an optimally modified PCS. Then
[…]
Proof.
First, assume that $\eta^{*}=\epsilon_{i_{0}}\text{ for some }i_{0}\in D_{1}$. As discussed in Proposition 1, we get $0\leq\eta_{bi_{0}},\eta_{i_{0}w},\eta_{bw}\leq\eta^{*}=\epsilon_{i_{0}}$ such that $\tilde{a}_{bi_{0}}^{*}\in\{a_{bi_{0}}-\eta_{bi_{0}},a_{bi_{0}}+\eta_{bi_{0}}\}$, $\tilde{a}_{i_{0}w}^{*}\in\{a_{i_{0}w}-\eta_{i_{0}w},a_{i_{0}w}+\eta_{i_{0}w}\}$ and $\tilde{a}_{bw}^{*}\in\{a_{bw}-\eta_{bw},a_{bw}+\eta_{bw}\}$. Since $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent, we have $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bw}^{*}$. Now, it is easy to observe that if any one of $\eta_{bi_{0}}$, $\eta_{i_{0}w}$ or $\eta_{bw}$ is strictly less than $\epsilon_{i_{0}}$, then at least one of the remaining two must be strictly greater than $\epsilon_{i_{0}}$, which is not possible. This gives $\eta_{bi_{0}}=\eta_{i_{0}w}=\eta_{bw}=\epsilon_{i_{0}}$. It is clear that the only combination among the eight possible values of $(\tilde{a}_{bi_{0}}^{*},\tilde{a}_{i_{0}w}^{*},\tilde{a}_{bw}^{*})$ that satisfies $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bw}^{*}$ is $(a_{bi_{0}}+\epsilon_{i_{0}},a_{i_{0}w}+\epsilon_{i_{0}},a_{bw}-\epsilon_{i_{0}})$. Therefore, $\tilde{a}_{bw}^{*}=a_{bw}-\epsilon_{i_{0}}$. The result follows by an analogous argument if $\eta^{*}=\epsilon_{j_{0}}\text{ for some }j_{0}\in D_{2}$.
Now, assume that $\eta^{*}=\epsilon_{i_{0},j_{0}}\text{ for some }i_{0},j_{0}\in D$. As discussed in Proposition 1, we get $0\leq\eta_{bi_{0}},\eta_{i_{0}w},\eta_{bj_{0}},\eta_{j_{0}w}\leq\eta^{*}=\epsilon_{i_{0},j_{0}}$ such that $\tilde{a}_{bi_{0}}^{*}\in\{a_{bi_{0}}-\eta_{bi_{0}},a_{bi_{0}}+\eta_{bi_{0}}\}$, $\tilde{a}_{i_{0}w}^{*}\in\{a_{i_{0}w}-\eta_{i_{0}w},a_{i_{0}w}+\eta_{i_{0}w}\}$, $\tilde{a}_{bj_{0}}^{*}\in\{a_{bj_{0}}-\eta_{bj_{0}},a_{bj_{0}}+\eta_{bj_{0}}\}$ and $\tilde{a}_{j_{0}w}^{*}\in\{a_{j_{0}w}-\eta_{j_{0}w},a_{j_{0}w}+\eta_{j_{0}w}\}$. Since $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent, we have $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bj_{0}}^{*}\times\tilde{a}_{j_{0}w}^{*}$. Now, it is easy to observe that if any one of $\eta_{bi_{0}}$, $\eta_{i_{0}w}$, $\eta_{bj_{0}}$ or $\eta_{j_{0}w}$ is strictly less than $\epsilon_{i_{0},j_{0}}$, then at least one of the remaining three must be strictly greater than $\epsilon_{i_{0},j_{0}}$, which is not possible. This gives $\eta_{bi_{0}}=\eta_{i_{0}w}=\eta_{bj_{0}}=\eta_{j_{0}w}=\epsilon_{i_{0},j_{0}}$. It is clear that the only combination among the sixteen possible values of $(\tilde{a}_{bi_{0}}^{*},\tilde{a}_{i_{0}w}^{*},\tilde{a}_{bj_{0}}^{*},\tilde{a}_{j_{0}w}^{*})$ that satisfies $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bj_{0}}^{*}\times\tilde{a}_{j_{0}w}^{*}$ is $(a_{bi_{0}}+\epsilon_{i_{0},j_{0}},a_{i_{0}w}+\epsilon_{i_{0},j_{0}},a_{bj_{0}}-\epsilon_{i_{0},j_{0}},a_{j_{0}w}-\epsilon_{i_{0},j_{0}})$. Therefore, $\tilde{a}_{bw}^{*}=\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. Hence the proof. ∎
Using equations (34) and (18a) in equation (20), we get optimal interval-weights.
Theorem 3.
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then the following statements hold.
  1. 1.
    If $\eta^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (25) is the only best optimally modified PCS.
  2. 2.
    If $\eta^{*}=\epsilon_{j_{0}}$ for some $j_{0}\in D_{2}$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (26) is the only best optimally modified PCS.
  3. 3.
    If $\eta^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$, then $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ given in equation (27) is the only best optimally modified PCS.
Proof.
Let $(\tilde{A}_{b}^{\prime*},\tilde{A}_{w}^{\prime*})$ be an optimally modified PCS. Let $|\tilde{a}_{bi}^{\prime*}-a_{bi}|=\eta_{bi}$ and $|\tilde{a}_{iw}^{\prime*}-a_{iw}|=\eta_{iw}$, where $i\in D$. So, $0\leq\eta_{bi},\eta_{iw}\leq\eta^{*}$. Also, $\tilde{a}_{bi}^{\prime*}\in\{a_{bi}+\eta_{bi},a_{bi}-\eta_{bi}\}$ and $\tilde{a}_{iw}^{\prime*}\in\{a_{iw}+\eta_{iw},a_{iw}-\eta_{iw}\}$.
First, assume that $\eta^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$. By Proposition 2, $\tilde{a}_{bw}^{*}=\tilde{a}_{bw}^{\prime*}=a_{bw}-\epsilon_{i_{0}}$. Therefore, $|\tilde{a}_{bw}^{*}-a_{bw}|=|\tilde{a}_{bw}^{\prime*}-a_{bw}|$.
Fix $i\in D$. If $a_{bi}\times a_{iw}\leq a_{bw}-\epsilon_{i_{0}}$, then by equation (28),
$|\tilde{a}_{bi}^{*}-a_{bi}|=|\tilde{a}_{iw}^{*}-a_{iw}|=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2},$
and thus,
$\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}.$
Since $(\tilde{A}_{b}^{\prime*},\tilde{A}_{w}^{\prime*})$ is consistent, we have $\tilde{a}_{bi}^{\prime}\times\tilde{a}_{iw}^{\prime}=\tilde{a}_{bw}^{\prime}=a_{bw}-\epsilon_{i_{0}}$. This, along with (29), implies that if one of $\eta_{bi}$ and $\eta_{iw}$ is strictly less than
$\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2},$
then the other necessarily exceeds it. This gives $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}\leq\max\{\eta_{bi},\eta_{iw}\}=\max\{|\tilde{a}_{bi}^{\prime*}-a_{bi}|,|\tilde{a}_{iw}^{\prime*}-a_{iw}|\}$. Moreover, if $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}=\max\{|\tilde{a}_{bi}^{\prime*}-a_{bi}|,|\tilde{a}_{iw}^{\prime*}-a_{iw}|\}$, then
$\eta_{bi}=\eta_{iw}=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2},$
which gives $\tilde{a}_{bi}^{\prime*}=\tilde{a}_{bi}^{*}$ and $\tilde{a}_{iw}^{\prime*}=\tilde{a}_{iw}^{*}$. Similar argument can be given if $a_{bi}\times a_{iw}>a_{bw}-\epsilon_{i_{0}}$. The conclusion also holds if $\eta^{*}=\epsilon_{j_{0}}\text{ for some }j_{0}\in D_{2}$ by an analogous reasoning.
Now, assume that $\eta^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$. By Proposition 2, $\tilde{a}_{bw}^{*}=\tilde{a}_{bw}^{\prime*}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. So, $|\tilde{a}_{bw}^{*}-a_{bw}|=|\tilde{a}_{bw}^{\prime*}-a_{bw}|$.
Fix $i\in D$. If $a_{bi}\times a_{iw}\leq(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$, then by equation (31),
$\begin{split}|\tilde{a}_{bi}-a_{bi}|&=|\tilde{a}_{iw}-a_{iw}|\\ &=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2},\end{split}$
and thus,
$\begin{split}\max&\{|\tilde{a}_{bi}-a_{bi}|,|\tilde{a}_{iw}-a_{iw}|\}\\ &=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}.\end{split}$
Since $(\tilde{A}_{b}^{\prime*},\tilde{A}_{w}^{\prime*})$ is consistent, we have $\tilde{a}_{bi}^{\prime}\times\tilde{a}_{iw}^{\prime}=\tilde{a}_{bw}^{\prime}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. This, along with (32), implies that if one of $\eta_{bi}$ and $\eta_{iw}$ is strictly less than
$\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2},$
then the other necessarily exceeds it. This gives $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}\leq\max\{\eta_{bi},\eta_{iw}\}=\max\{|\tilde{a}_{bi}^{\prime*}-a_{bi}|,|\tilde{a}_{iw}^{\prime*}-a_{iw}|\}$. Moreover, if $\max\{|\tilde{a}_{bi}^{*}-a_{bi}|,|\tilde{a}_{iw}^{*}-a_{iw}|\}=\max\{|\tilde{a}_{bi}^{\prime*}-a_{bi}|,|\tilde{a}_{iw}^{\prime*}-a_{iw}|\}$, then
$\eta_{bi}=\eta_{iw}=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2},$
which gives $\tilde{a}_{bi}^{\prime*}=\tilde{a}_{bi}^{*}$ and $\tilde{a}_{iw}^{\prime*}=\tilde{a}_{iw}^{*}$. Similar argument can be given if $a_{bi}\times a_{iw}>(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})$. Hence the proof. ∎
Using the best optimally modified PCS, the best optimal weight set is determined via equation (1), yielding the final resultant weights.
4.1.2 Calculation of Weights in the Presence of Multiple Best or Worst Criteria
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of decision criteria with $c_{b_{1}},c_{b_{2}},\ldots,c_{b_{n_{1}}}$ as the best and $c_{w_{1}},c_{w_{2}},\ldots,$ $c_{w_{n_{2}}}$ as the worst criteria, and let $(A_{b},A_{w})$ be a PCS. Therefore,
$a_{b_{1}i}=a_{b_{2}i}=\ldots=a_{b_{n_{1}}i}=a_{bi}\text{ (say)},\quad a_{iw_{1}}=a_{iw_{2}}=\ldots=a_{iw_{n_{2}}}=a_{iw}\text{ (say)},$
$a_{b_{1}w_{1}}=\ldots a_{b_{1}w_{n_{2}}}=a_{b_{2}w_{1}}=\ldots a_{b_{2}w_{n_{2}}}=\ldots=a_{b_{n_{1}}w_{1}}=\ldots a_{b_{n_{1}}w_{n_{2}}}=a_{bi}\text{ (say)}$
for $i\in D^{\prime}$, where $D^{\prime}=C\setminus\{c_{b_{1}},\ldots,c_{b_{n_{1}}},c_{w_{1}}\ldots,c_{w_{n_{2}}}\}$. Whenever there is no ambiguity, we simply use the notation $D^{\prime}=\{1,\ldots,n\}\setminus\{b_{1},\ldots,b_{n_{1}},w_{1},\ldots,w_{n_{2}}\}$.
To ensure no weight difference for criteria having equal preference, instead of considering the system of equations (1), we consider the system of equations
$\begin{split}&\frac{w_{b_{1}}}{w_{i}}=\frac{w_{b_{2}}}{w_{i}}=\ldots=\frac{w_{b_{n_{1}}}}{w_{i}}=a_{bi},\quad\frac{w_{i}}{w_{w_{1}}}=\frac{w_{i}}{w_{w_{2}}}=\ldots=\frac{w_{i}}{w_{w_{n_{2}}}}=a_{iw},\\ &\frac{w_{b_{1}}}{w_{w_{1}}}=\ldots=\frac{w_{b_{1}}}{w_{w_{n_{2}}}}=\frac{w_{b_{2}}}{w_{w_{1}}}=\ldots=\frac{w_{b_{2}}}{w_{w_{n_{2}}}}=\ldots=\frac{w_{b_{n_{1}}}}{w_{w_{1}}}=\ldots=\frac{w_{b_{n_{1}}}}{w_{w_{n_{2}}}}=a_{bw},\quad i\in D^{\prime},\\ &w_{1}+w_{2}+\ldots+w_{n}=1,\end{split}$ (35)
which is equivalent to the system of equations
$\begin{split}&\frac{w_{b_{1}}}{w_{i}}=a_{bi},\quad\frac{w_{i}}{w_{w_{1}}}=a_{iw},\quad\frac{w_{b_{1}}}{w_{w_{1}}}=a_{bw},\quad i\in D^{\prime},\\ &w_{b_{1}}=w_{b_{2}}=\ldots=w_{b_{n_{1}}},\quad w_{w_{1}}=w_{w_{2}}=\ldots=w_{w_{n_{2}}},\quad w_{1}+w_{2}+\ldots+w_{n}=1.\end{split}$ (36)
Consider the following minimization problem.
$\begin{split}&\min\epsilon\\ &\text{subject to:}\\ &\left|\frac{w_{b_{1}}}{w_{i}}-a_{bi}\right|\leq\epsilon,\quad\left|\frac{w_{i}}{w_{w_{1}}}-a_{iw}\right|\leq\epsilon,\quad\left|\frac{w_{b_{1}}}{w_{w_{1}}}-a_{bw}\right|\leq\epsilon,\\ &n_{1}w_{b_{1}}+n_{2}w_{w_{1}}+\sum_{k\in D^{\prime}}w_{k}=1,\quad w_{j}\geq 0\text{ for all }i\in D^{\prime}\text{ and }j\in C.\end{split}$ (37)
Each optimal solution of problem (37), along with $w_{b_{1}}^{*}=w_{b_{2}}^{*}=\ldots=w_{b_{n_{1}}}^{*}$ and $w_{w_{1}}^{*}=w_{w_{2}}^{*}=\ldots=w_{w_{n_{2}}}^{*}$, gives $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$ is an optimal weight set and $\epsilon^{*}$ is a measurement of its accuracy. Since $\epsilon^{*}$ is also the optimal objective value, it remains the same for all $W^{*}$.
When problem (37) has multiple optimal solutions, the optimal interval-weight for criterion $c_{k}$ is $w_{k}^{\prime*},w_{k}^{\prime\prime*}]$, where $w_{k}^{\prime*}$ and $w_{k}^{\prime\prime*}$ represent the optimal objective values of problems
$\begin{split}&\min w_{k}\\ &\text{subject to:}\\ &\left|\frac{w_{b_{1}}}{w_{i}}-a_{bi}\right|\leq\epsilon^{*},\quad\left|\frac{w_{i}}{w_{w_{1}}}-a_{iw}\right|\leq\epsilon^{*},\quad\left|\frac{w_{b_{1}}}{w_{w_{1}}}-a_{bw}\right|\leq\epsilon^{*},\\ &n_{1}w_{b_{1}}+n_{2}w_{w_{1}}+\sum_{k\in D^{\prime}}w_{k}=1,\quad w_{j}\geq 0\text{ for all }i\in D^{\prime}\text{ and }j\in C\end{split}$ (38)
and
$\begin{split}&\max w_{k}\\ &\text{subject to:}\\ &\left|\frac{w_{b_{1}}}{w_{i}}-a_{bi}\right|\leq\epsilon^{*},\quad\left|\frac{w_{i}}{w_{w_{1}}}-a_{iw}\right|\leq\epsilon^{*},\quad\left|\frac{w_{b_{1}}}{w_{w_{1}}}-a_{bw}\right|\leq\epsilon^{*},\\ &n_{1}w_{b_{1}}+n_{2}w_{w_{1}}+\sum_{k\in D^{\prime}}w_{k}=1,\quad w_{j}\geq 0\text{ for all }i\in D^{\prime}\text{ and }j\in C\end{split}$ (39)
respectively.
Consider the following minimization problem.
$\begin{split}&\min\eta\\ &\text{subject to:}\\ &\left|\tilde{a}_{bi}-a_{bi}\right|\leq\eta,\quad\left|\tilde{a}_{iw}-a_{iw}\right|\leq\eta,\quad\left|\tilde{a}_{bw}-a_{bw}\right|\leq\eta,\\ &\tilde{a}_{bi}\times\tilde{a}_{iw}=\tilde{a}_{bw},\quad\tilde{a}_{bi},\tilde{a}_{iw},\tilde{a}_{bw}\geq 0\text{ for all }i\in D^{\prime}.\end{split}$ (40)
Each optimal solution of problem (40), along with $\tilde{a}_{bb}^{*}=\tilde{a}_{ww}^{*}=1$, leads to an optimally modified PCS $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ and $\eta^{*}$ is a measurement of its accuracy. Since $\eta^{*}$ is also the optimal objective value, it remains the same for all $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$.
Analogous to the case with unique best and worst criteria, it can be verified that equation (8) holds for the collections of optimal solutions of problems (37) and (40), and thus, the two problems are equivalent.
Theorem 4.
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\epsilon^{*}$ and $\eta^{*}$ be the optimal objective value of problems (37) and (40). Then
$\epsilon^{*}=\eta^{*}=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}.$ (41)
Also, the collection of all optimally modified PCS is
[…]
Proof.
The result can be proven by replicating the arguments of Theorem 2 and Proposition 2 and is thus omitted. ∎
Using equation (42) in equation (19), we get
$\begin{split}{w_{b_{1}}^{l}}^{*}&={w_{b_{2}}^{l}}^{*}=\ldots={w_{b_{n_{1}}}^{l}}^{*}=\frac{\tilde{a}_{bw}^{*}}{n_{2}+n_{1}\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D^{\prime}}\min\biggl{\{}a_{jw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}-\epsilon^{*}}\biggr{\}}},\\ {w_{b_{1}}^{u}}^{*}&={w_{b_{2}}^{u}}^{*}=\ldots={w_{b_{n_{1}}}^{u}}^{*}=\frac{\tilde{a}_{bw}^{*}}{n_{2}+n_{1}\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D^{\prime}}\max\biggl{\{}a_{jw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}+\epsilon^{*}}\biggr{\}}},\\ {w_{w_{1}}^{l}}^{*}&={w_{w_{2}}^{l}}^{*}=\ldots={w_{w_{n_{2}}}^{l}}^{*}=\frac{1}{n_{2}+n_{1}\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D^{\prime}}\min\biggl{\{}a_{jw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}-\epsilon^{*}}\biggr{\}}},\\ {w_{w_{1}}^{u}}^{*}&={w_{w_{2}}^{u}}^{*}=\ldots={w_{w_{n_{2}}}^{u}}^{*}=\frac{1}{n_{2}+n_{1}\tilde{a}_{bw}^{*}+\displaystyle\sum_{j\in D^{\prime}}\max\biggl{\{}a_{jw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}+\epsilon^{*}}\biggr{\}}},\\ {w_{i}^{l}}^{*}&=\frac{\max\biggl{\{}a_{iw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}+\epsilon^{*}}\biggr{\}}}{n_{2}+n_{1}\tilde{a}_{bw}^{*}+\max\biggl{\{}a_{iw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}+\epsilon^{*}}\biggr{\}}+\displaystyle\sum_{\begin{subarray}{c}j\in D^{\prime}\\ j\neq i\end{subarray}}\min\biggl{\{}a_{jw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}-\epsilon^{*}}\biggr{\}}},\\ {w_{i}^{u}}^{*}&=\frac{\min\biggl{\{}a_{iw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}-\epsilon^{*}}\biggr{\}}}{n_{2}+n_{1}\tilde{a}_{bw}^{*}+\min\biggl{\{}a_{iw}+\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bi}-\epsilon^{*}}\biggr{\}}+\displaystyle\sum_{\begin{subarray}{c}j\in D^{\prime}\\ j\neq i\end{subarray}}\max\biggl{\{}a_{jw}-\epsilon^{*},\frac{\tilde{a}_{bw}^{*}}{a_{bj}+\epsilon^{*}}\biggr{\}}},\quad\text{where }i\in D^{\prime}.\end{split}$ (43)
Theorem 5.
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\epsilon^{*}$ and $\eta^{*}$ be the optimal objective value of problems (37) and (40). Then the following statements hold.
  1. 1.
    If $\eta^{*}=\epsilon_{i_{0}}$ for some $i_{0}\in D_{1}$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
    $\begin{split}&\tilde{a}_{bi}=\frac{a_{bi}-a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2},\\ &\tilde{a}_{iw}=\frac{-a_{bi}+a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\quad\text{for all }i\in D^{\prime},\\ &\tilde{a}_{bw}=a_{bw}-\epsilon_{i_{0}}\end{split}$ (44)
    is the only best optimally modified PCS.
  2. 2.
    If $\eta^{*}=\epsilon_{j_{0}}$ for some $j_{0}\in D_{2}$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
    $\begin{split}&\tilde{a}_{bi}=\frac{a_{bi}-a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}+\epsilon_{j_{0}})}}{2},\\ &\tilde{a}_{iw}=\frac{-a_{bi}+a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}+\epsilon_{j_{0}})}}{2}\quad\text{for all }i\in D^{\prime},\\ &\tilde{a}_{bw}=a_{bw}+\epsilon_{j_{0}}\end{split}$ (45)
    is the only best optimally modified PCS.
  3. 3.
    If $\eta^{*}=\epsilon_{i_{0},j_{0}}$ for some $i_{0},j_{0}\in D$, then $(\tilde{A}_{b},\tilde{A}_{w})$ defined as
    $\begin{split}&\tilde{a}_{bi}=\frac{a_{bi}-a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2},\\ &\tilde{a}_{iw}=\frac{-a_{bi}+a_{iw}+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{for all }i\in D^{\prime},\\ &\tilde{a}_{bw}=(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})\end{split}$ (46)
    is the only best optimally modified PCS.
Using the best optimally modified PCS, the best optimal weight set is determined via equation (36), yielding the final resultant weights.
4.2 Consistency Analysis
In this subsection, our aim is to derive analytical expressions for CI and CR. From equation (33), it follows that CI can be obtained by finding the maximum possible values of $\epsilon_{i}$ and $\epsilon_{i,j}$.
Proposition 3.
Let $i,j\in D_{1}$ be such that $a_{bi}\leq a_{bj}$ and $a_{iw}\leq a_{jw}$. Then $\epsilon_{j}\leq\epsilon_{i}$.
Proof.
Suppose, if possible, $\epsilon_{i}<\epsilon_{j}$. Then we get $a_{bi}+\epsilon_{i}<a_{bj}+\epsilon_{j}$ and $a_{iw}+\epsilon_{i}<a_{jw}+\epsilon_{j}$. This gives
$(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})<(a_{bj}+\epsilon_{j})\times(a_{jw}+\epsilon_{j}).$ (47)
From equation (10), we have $(a_{bi}+\epsilon_{i})\times(a_{iw}+\epsilon_{i})=a_{bw}-\epsilon_{i}$ and $(a_{bj}+\epsilon_{j})\times(a_{jw}+\epsilon_{j})=a_{bw}-\epsilon_{j}$. Using these in equation (47), we get $a_{bw}-\epsilon_{i}<a_{bw}-\epsilon_{j}$, which is contradiction as $\epsilon_{i}<\epsilon_{j}$ implies $a_{bw}-\epsilon_{j}<a_{bw}-\epsilon_{i}$. Thus, $\epsilon_{j}\leq\epsilon_{i}$. Hence the proof. ∎
Proposition 4.
Let $i,j\in D_{2}$ be such that $a_{bi}\leq a_{bj}$ and $a_{iw}\leq a_{jw}$. Then $\epsilon_{i}\leq\epsilon_{j}$.
Proposition 5.
Let $i,j,k\in D$ be such that $a_{bi}\leq a_{bj}\leq a_{bk}$ and $a_{iw}\leq a_{jw}\leq a_{kw}$. Then $\epsilon_{i,j}\leq\epsilon_{i,k}$ and $\epsilon_{j,k}\leq\epsilon_{i,k}$.
The proofs of Proposition 4 and Proposition 5 are similar to the proof of Proposition 3 and thus omitted.
Proposition 3 implies that for any scale satisfying $a_{bi},a_{iw}\geq 1$ for all $i\in C$, the maximum possible value of $\epsilon_{i}$, where $i\in D_{1}$, is attained when $a_{bi}=a_{iw}=1$ for some $i\in D_{1}$. So,
$\max\{\epsilon_{i}:i\in D_{1}\}=\frac{-3+\sqrt{4a_{bw}+5}}{2}.$
Proposition 4 shows that for any scale satisfying $a_{bw}\geq 1$, the maximum possible value of $\epsilon_{i}$, where $i\in D_{2}$, is attained when $a_{bi}=a_{iw}=a_{bw}$ for some $i\in D_{2}$. So,
$\max\{\epsilon_{i}:i\in D_{2}\}=\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}.$
Proposition 5 suggests that for any scale satisfying $a_{bi},a_{iw}\geq 1$ for all $i\in C$, the maximum possible value of $\epsilon_{i,j}$ is attained when $a_{bi}=a_{iw}=1$ for some $i\in D_{1}$ and $a_{bj}=a_{jw}=a_{bw}$ for some $j\in D_{2}$. So,
$\max\{\epsilon_{i,j}:i,j\in D\}=\frac{a_{bw}^{2}-1}{2a_{bw}+2}.$
From the above discussion, it follows that
$\text{CI}_{a_{bw}}=\max\biggl{\{}\frac{-3+\sqrt{4a_{bw}+5}}{2},\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr{\}}.$
Let
$f(x)=\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}-\frac{-3+\sqrt{4a_{bw}+5}}{2},\ x\in[1,\infty).$
Note that
$f^{\prime}(x)=1-\frac{2}{\sqrt{8a_{bw}+1}}-\frac{1}{\sqrt{4a_{bw}+5}}.$
Since $a_{bw}\geq 1$, we get $f^{\prime}(x)\geq 0$, i.e., $f$ is increasing. Therefore, $f(x)\geq f(1)=0$ for all $x\in[1,\infty).$ This gives
$\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2}\geq\frac{-3+\sqrt{4a_{bw}+5}}{2},$
and thus,
$\text{CI}_{a_{bw}}=\max\biggl{\{}\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr{\}},$ (48)
which is analytical expression for CI. The values of CI for some established scales are given in Table 9.
Table 9: The values of CI for some scales
Saaty Salo-Hämäläinen Lootsma Donegan-Dodd-
scale scale scale McMaster scale
$a_{bw}$ CI $a_{bw}$ CI $a_{bw}$ CI $a_{bw}$ CI
22 $0.5$ $1.2222$ $0.1111$ $\sqrt{2}$ $0.2071$ $1.1257$ $0.0629$
33 11 $1.5$ $0.25$ 22 $0.5$ $1.2715$ $0.1358$
44 $1.6277$ $1.8571$ $0.4286$ $2\sqrt{2}$ $0.9142$ $1.4470$ $0.2235$
55 $2.2984$ $2.3333$ $0.6667$ 44 $1.6277$ $1.6684$ $0.3342$
66 33 33 11 $4\sqrt{2}$ $2.7563$ $1.9670$ $0.4835$
77 $3.7250$ 44 $1.6277$ 88 $4.4688$ $2.4142$ $0.7071$
88 $4.4688$ $5.6667$ $2.7633$ $8\sqrt{2}$ $7.0307$ $3.2289$ $1.1390$
99 $5.2279$ 99 $5.2279$ $16$ $10.8211$ $5.8284$ $2.8778$
Now, substituting the expressions for $\epsilon^{*}$ and CI from equations (33) and (48), respectively, into equation (6), we get
$\text{CR}=\frac{\max\left\{\bigg{|}\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg{|},\bigg{|}\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}\bigg{|}:i,j\in D\right\}}{\max\biggl{\{}\frac{2a_{bw}+1-\sqrt{8a_{bw}+1}}{2},\frac{a_{bw}^{2}-1}{2a_{bw}+2}\biggr{\}}},$ (49)
which is analytical expression for CR.
Some important points related to the analytical expressions for CI and CR are as follows.
  1. 1.
[…]
  2. 2.
    A consistency indicator that provides instant feedback to DMs regarding judgment inconsistencies improves the effectiveness of an MCDM method [26]. Equation (6) represents an output-based formulation of CR that lacks this ability and can only provide feedback after completing the entire calculation process. This limitation is addressed by the analytical expression of CR given in equation (49), which represents an input-based formulation of CR.
  3. 3.
    For a consistency indicator to demonstrate reasonable behavior, it must satisfy some specific properties [7, 23]. Proposition 6 outlines several of these properties without proof, as their proofs are analogous to those in [26, Proposition 1].
Proposition 6.
CR exhibits the following properties.
  1. 1.
    CR is normalized, i.e., $0\leq$ CR $\leq 1$.
  2. 2.
    CR =0=0 if and only if $(A_{b},A_{w})$ is consistent.
  3. 3.
    CR exhibits permutation invariance with respect to the criteria indices.
  4. 4.
    CR is non-increasing with respect to criterion elimination.
  5. 5.
    CR is a continuous function of $a_{bi},a_{iw}$ and $a_{bw}$.
  6. 6.
    For a consistent $(A_{b},A_{w})$, CR increases when either $a_{bi}$ or $a_{iw}$ moves away from its original value in the range $1,a_{bw}]$.
4.3 Numerical Examples
In this subsection, we revisit the six examples (Example 1–Example 6) from the section “Research Gap” to demonstrate and validate the proposed approach.
Example 1: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.4056$, $\epsilon_{3}=0.3944$, $\epsilon_{4}=0.5363$, $\epsilon_{4,2}=0.5163$ and $\epsilon_{4,3}=0.5422$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=0.5422$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=8.4987$. From equation (20), we get
$\begin{split}{w_{1}^{l}}^{*}&=\frac{\tilde{a}_{15}^{*}}{1+\tilde{a}_{15}^{*}+\displaystyle\sum_{j=2,3,4}\min\biggl{\{}a_{j5}+\epsilon^{*},\frac{\tilde{a}_{15}^{*}}{a_{1j}-\epsilon^{*}}\biggr{\}}}\\ &=\frac{8.4987}{1+8.4987+1.0048+3.4578+3.5422}\\ &=0.4855.\end{split}$
Similarly, ${w_{1}^{u}}^{*}=0.4868$. Therefore, $w_{1}^{*}=[0.4855,0.4868]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0549,0.0574]$, $w_{3}^{*}=[0.1975,0.1981]$, $w_{4}^{*}=[0.2024,0.2029]$ and $w_{5}^{*}=[0.0571,0.0573]$.
Since $\epsilon^{*}=\epsilon_{4,3}$, by statement 3 of Theorem 3,
$\begin{split}&\tilde{a}_{bi}^{*}=\frac{a_{1i}-a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})}}{2},\\ &\tilde{a}_{iw}^{*}=\frac{-a_{1i}+a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})}}{2},\end{split}$
where $i=2,3,4$, along with $\tilde{a}_{11}^{*}=\tilde{a}_{55}^{*}=1$ and $\tilde{a}_{15}^{*}=8.4987$, form the best optimally modified PCS. Therefore,
$\tilde{A}_{b}^{*}=(1,8.4962,2.4519,2.3934,8.4638),\quad\tilde{A}_{w}^{*}=(8.4638,0.9962,3.4519,3.5363,1)^{T}$
[…]
Table 3 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 2: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=1.6054$, $\epsilon_{3}=1.4764$, $\epsilon_{4}=1.7228$, $\epsilon_{4,2}=1.7882$ and $\epsilon_{4,3}=1.7999$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=1.7999$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=14.8760$. From equation (20), we get
$\begin{split}{w_{1}^{l}}^{*}&=\frac{\tilde{a}_{15}^{*}}{1+\tilde{a}_{15}^{*}+\displaystyle\sum_{j=2,3,4}\min\biggl{\{}a_{j5}+\epsilon^{*},\frac{\tilde{a}_{15}^{*}}{a_{1j}-\epsilon^{*}}\biggr{\}}}\\ &=\frac{14.8760}{1+14.8760+1.0476+3.8569+3.2141}\\ &=0.6200.\end{split}$
Similarly, ${w_{1}^{u}}^{*}=0.6205$. Therefore, $w_{1}^{*}=[0.6200,0.6205]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0429,0.0437]$, $w_{3}^{*}=[0.1607,0.1609]$, $w_{4}^{*}=[0.1340,0.1341]$ and $w_{5}^{*}=[0.0416,0.0417]$.
Since $\epsilon^{*}=\epsilon_{4,3}$, by statement 3 of Theorem 3,
$\begin{split}&\tilde{a}_{bi}^{*}=\frac{a_{1i}-a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})}}{2},\\ &\tilde{a}_{iw}^{*}=\frac{-a_{1i}+a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})}}{2},\end{split}$
where $i=2,3,4$, along with $\tilde{a}_{11}^{*}=\tilde{a}_{55}^{*}=1$ and $\tilde{a}_{15}^{*}=14.8760$, form the best optimally modified PCS. Therefore,
$\tilde{A}_{b}^{*}=(1,14.2179,3.8569,4.6283,14.8760),\quad\tilde{A}_{w}^{*}=(14.8760,1.0463,3.8569,3.2141,1)^{T}$
[…]
Table 4 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 3: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.6959$, $\epsilon_{3}=0.6781$, $\epsilon_{4}=0.8086$, $\epsilon_{4,2}=0.8825$ and $\epsilon_{4,3}=0.8975$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=0.8975$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=5.4354$. From equation (20), we get
$\begin{split}{w_{1}^{l}}^{*}&=\frac{\tilde{a}_{15}^{*}}{1+\tilde{a}_{15}^{*}+\displaystyle\sum_{j=2,3,4}\min\biggl{\{}a_{j5}+\epsilon^{*},\frac{\tilde{a}_{15}^{*}}{a_{1j}-\epsilon^{*}}\biggr{\}}}\\ &=\frac{5.4354}{1+5.4354+1.1023+2.3314+2.8645}\\ &=0.4269.\end{split}$
Similarly, ${w_{1}^{u}}^{*}=0.4280$. Therefore, $w_{1}^{*}=[0.4269,0.4280]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0842,0.0866]$, $w_{3}^{*}=[0.1831,0.1836]$, $w_{4}^{*}=[0.2250,0.2255]$ and $w_{5}^{*}=[0.0785,0.0787]$.
Since $\epsilon^{*}=\epsilon_{4,3}$, by statement 3 of Theorem 3,
$\begin{split}&\tilde{a}_{bi}^{*}=\frac{a_{1i}-a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})}}{2},\\ &\tilde{a}_{iw}^{*}=\frac{-a_{1i}+a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})}}{2},\end{split}$
where $i=2,3,4$, along with $\tilde{a}_{11}^{*}=\tilde{a}_{55}^{*}=1$ and $\tilde{a}_{15}^{*}=5.4354$, form the best optimally modified PCS. Therefore,
$\tilde{A}_{b}^{*}=(1,4.9577,2.3314,1.8975,5.4354),\quad\tilde{A}_{w}^{*}=(5.4354,1.0963,2.3314,2.8645,1)^{T}$
[…]
Table 5 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 4: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.4355$, $\epsilon_{3}=0.4401$, $\epsilon_{4}=0.4865$, $\epsilon_{4,2}=0.0.5480$ and $\epsilon_{4,3}=0.5458$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,2}=0.5480$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,2})\times(a_{45}+\epsilon_{4,2})=6.8344$. From equation (20), we get
$\begin{split}{w_{1}^{l}}^{*}&=\frac{\tilde{a}_{15}^{*}}{1+\tilde{a}_{15}^{*}+\displaystyle\sum_{j=2,3,4}\min\biggl{\{}a_{j5}+\epsilon^{*},\frac{\tilde{a}_{15}^{*}}{a_{1j}-\epsilon^{*}}\biggr{\}}}\\ &=\frac{6.8344}{1+6.8344+2.6143+4.0487+2.2801}\\ &=0.4074.\end{split}$
Similarly, ${w_{1}^{u}}^{*}=0.4077$. Therefore, $w_{1}^{*}=[0.4074,0.4077]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.1558,0.1560]$, $w_{3}^{*}=[0.2407,0.2413]$, $w_{4}^{*}=[0.1359,0.1360]$ and $w_{5}^{*}=[0.0596,0.0597]$.
Since $\epsilon^{*}=\epsilon_{4,2}$, by statement 3 of Theorem 3,
$\begin{split}&\tilde{a}_{bi}^{*}=\frac{a_{1i}-a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,2})\times(a_{45}+\epsilon_{4,2})}}{2},\\ &\tilde{a}_{iw}^{*}=\frac{-a_{1i}+a_{i5}+\sqrt{(a_{1i}+a_{i5})^{2}-4\times a_{1i}\times a_{i5}+4(a_{14}+\epsilon_{4,2})\times(a_{45}+\epsilon_{4,2})}}{2},\end{split}$
where $i=2,3,4$, along with $\tilde{a}_{11}^{*}=\tilde{a}_{55}^{*}=1$ and $\tilde{a}_{15}^{*}=6.8344$, form the best optimally modified PCS. Therefore,
$\tilde{A}_{b}^{*}=(1,2.6143,1.6923,2.9975,6.8344),\quad\tilde{A}_{w}^{*}=(6.8344,2.6143,4.0388,2.2801,1)^{T}$
[…]
Table 6 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 5: Here, $D_{1}=\{2\}$, $D_{2}=\{3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.3028$, $\epsilon_{3}=0.4384$ and $\epsilon_{2,3}=0.5$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{2,3}\}=\epsilon_{2,3}=0.5$. Now, equation 34 gives $\tilde{a}_{14}^{*}=(a_{12}+\epsilon_{2,3})\times(a_{24}+\epsilon_{2,3})=2.25$. From equation (20), we get
$\begin{split}{w_{1}^{l}}^{*}&=\frac{\tilde{a}_{14}^{*}}{1+\tilde{a}_{14}^{*}+\displaystyle\sum_{j=2,3}\min\biggl{\{}a_{j4}+\epsilon^{*},\frac{\tilde{a}_{14}^{*}}{a_{1j}-\epsilon^{*}}\biggr{\}}}\\ &=\frac{2.25}{1+2.25+1.5+1.5}\\ &=0.36.\end{split}$
[…]
Table 7 shows that the obtained $\epsilon^{*}$ and the optimal weight set coincide with their actual values, validating the proposed framework.
Example 6: Here, $n_{1}=2$, $n_{2}=1$, $D_{1}=\{3\}$ and $D_{2}=D_{3}=\phi$. By equations (13), we get $\epsilon_{2}=0.1623$. So, from equation (41), it follows that $\epsilon^{*}=\epsilon_{2}=0.1623$. Now, equation 42 gives $\tilde{a}_{14}^{*}=\tilde{a}_{24}^{*}=(a_{13}+\epsilon_{2})\times(a_{34}+\epsilon_{2})=6.8378$. From equation (43), we get
$\begin{split}{w_{1}^{l}}^{*}={w_{2}^{l}}^{*}&=\frac{\tilde{a}_{14}^{*}}{1+2\tilde{a}_{14}^{*}+\min\biggl{\{}a_{34}+\epsilon^{*},\frac{\tilde{a}_{14}^{*}}{a_{13}-\epsilon^{*}}\biggr{\}}}\\ &=\frac{6.8378}{1+2\times 6.8378+3.1623}\\ &=0.3833.\end{split}$
[…]
Note that both $c_{1}$ and $c_{2}$ have the same optimal weights.
5 Conclusions and Future Directions
The BWM has emerged as a powerful MCDM tool, widely adopted in real-world applications. This study introduces a generalized analytical framework for its nonlinear model, compatible with any scale and any number of DMs. The framework derives closed-form solutions for optimal interval-weights, CI and CR, while also identifying the best weight set from all possible solutions. In the presence of multiple best/worst criteria, a key modification in the original optimization model ensures the selection of optimal weight sets in which criteria with equal preferences receive the same weights. The research advances BWM methodology by enhancing its capability to handle both unique and multiple optimal solutions across all preference values. By eliminating the need for specialized optimization software, the framework significantly improves computational efficiency. It also rectifies previously reported inaccuracies in CI calculations and transforms CR into an effective real-time consistency measure for immediate DM feedback.
Several important challenges remain unresolved in this research. Specifically, analytical solutions for optimal weights have not yet been developed for certain BWM variants, including the Euclidean BWM [22] and $\alpha$-FBWM [30].
Acknowledgements
The first author gratefully acknowledges the Council of Scientific & Industrial Research (CSIR), India for financial support to carry out the research work.
Declaration of Conflict of Interest
The authors declare that they have no known conflict of financial interests or personal relationships that could have appeared to influence the work reported in this paper.